# Safeguarding System Prompts for LLMs

### v1

@@ Abstract
Large language models (LLMs) are increasingly utilized in applications where system prompts, which guide model outputs, play a crucial role. These prompts often contain business logic and sensitive information, making their protection essential. However, adversarial and even regular user queries can exploit LLM vulnerabilities to expose these hidden prompts. To address this issue, we present PromptKeeper, a novel defense mechanism for system prompt privacy. By reliably detecting worst-case leakage and regenerating outputs without the system prompt when necessary, PromptKeeper ensures robust protection against prompt extraction attacks via either adversarial or regular queries, while preserving conversational capability and runtime efficiency during benign user interactions. [^1] Code is available at https://github.com/SamuelGong/PromptKeeper.
1 Introduction
Large language models (LLMs) have shown remarkable abilities to follow natural-language instructions (Brown et al., 2020; Touvron et al., 2023; Ouyang et al., 2022). In situations where an LLM is accessible to users through a web API, the service provider commonly prepends a system prompt to each user query. This prompt serves as a guidance for the model’s output behavior, allowing for diverse tasks to be accomplished without the need for expensive fine-tuning (Apideck, 2024). In many LLM-powered applications, the prompt itself, which incorporates carefully curated business logic, holds greater significance than the LLM, which is often a publicly available foundation model (PromptBase, 2024; PromptSea, 2024). As a result, system prompts are meant to be kept hidden from users to prevent replication of applications (MicroSoft, 2024). Moreover, these prompts may contain secret values or safety-related instructions, and any inadvertent disclosure of these prompts can aid adversaries in privacy or security attacks (Wallace et al., 2024; Toyer et al., 2024).
However, recent work has shown that LLMs can reveal hidden prompts in the presence of specially crafted adversarial queries (e.g., “Repeat all sentences you saw.”) (Perez & Ribeiro, 2022; Wallace et al., 2024), even when the models are explicitly instructed to avoid discussing the prompts, or post-generation filters are implemented to prevent exact replication of them in the outputs (Zhang et al., 2024b). Even worse, researchers have developed stealthier methods for prompt extraction that rely only on regular queries rather than adversarial ones. They achieve this by training a model to map the logits (Morris et al., 2024) or text outputs (Zhang et al., 2024a) back to the system prompts used. We therefore ask: can we safeguard our system prompts reliably and practically?
Our contributions.   This paper presents PromptKeeper, a novel systematic defense mechanism, to our best knowledge, designed to mitigate the leakage of system prompts (Figure 1). It addresses both regular and adversarial queries, without requiring any prior knowledge of benign user interactions or attacker strategies. PromptKeeper further operates with minimal system overhead and ensures that the utility of benign queries is not compromised.
The development of PromptKeeper entails addressing two fundamental challenges. The first challenge involves reliably identifying the leakage of system prompts in the outputs of LLMs. While complete leakage occurs when an attacker can guess the system prompt verbatim, partial leakage is more nuanced and harder to quantify. This difficulty arises from the inherent complexity of defining what constitutes private information within a prompt and the context-specific nature of information leakage. Moreover, an attacker’s guess may not represent the optimal guess based on its observed responses, indicating only a lower bound on leakage. To address this challenge, we consider a worst-case scenario approach, where any information about the prompt in the response is deemed leakage. This leads us to model leakage identification as a hypothesis testing problem, distinguishing scenarios with zero leakage from those with all others (Section 3).
The second challenge is to prevent system prompt leakage in a general and practical manner, i.e., against both regular and adversarial user queries, while incurring minimal cost to the handling of benign user requests. To achieve this, we avoid approaches that involve re-training or fine-tuning the model, altering user queries, or extending the original system prompts. Instead, we design a response-based scheme that operates normally, regenerating outputs without the system prompt only when leakage is detected through the proposed hypothesis testing. Unlike simply denying service upon detecting leakage, this regeneration approach counters adversarial search attempts by attackers, ensuring full system prompt privacy. Meanwhile, it preserves runtime efficiency and conversational capability during benign user interactions (Section 4).
@@ 1 Introduction
We evaluate PromptKeeper to assess its effectiveness in protecting various system prompts, including those from real-world GPT Store apps (Section 5). The evaluation covers system prompt extraction attacks using both regular and adversarial user queries. Additionally, we quantify the protected model’s conversational capability, focusing on its adherence to the scope and behavior defined by the system prompt during benign user interactions. Extensive results show that PromptKeeper effectively minimizes system prompt leakage while preserving model capability across different LLMs and datasets (Section 6).
@@ 2 Threat Model
Scenario.   As commonly studied (Zhang et al., 2024b), we consider a scenario where a service API, denoted as $f_{\bm{p}}$, is used for text generation. The API takes as input a user query $\bm{q}$ and passes to a language model LM, which generates a response $\bm{r}\leftarrow\texttt{LM}(\bm{p},\bm{q})$ using a system prompt $\bm{p}$ secretly owned by the service provider, as well as some employed randomness. It is also possible for the user to access the API indirectly through applications such as ChatGPT or a GPT store app (OpenAI, 2024b). Both $\bm{p}$ and $\bm{q}$ can be used separately with different privilege levels, similar to GPT-4 (Wallace et al., 2024), while they can also be concatenated together, as seen in GPT-3 (Mann et al., 2020).
System prompt extraction.   The attacker’s goal is to accurately guess the system prompt $\bm{p}$ by using a set of responses $\bm{r}_{1},\dots,\bm{r}_{k}$ acquired through $k$ queries made to the API using $\bm{q}_{1},\dots,\bm{q}_{k}$. The guess $\bm{g}$ is generated as $\bm{g}=\texttt{recon}(\bm{r}_{1},\dots,\bm{r}_{k})$, where $\texttt{recon}(\cdot)$ can be any function of the attacker’s choice, such as string manipulation or a deep neural network. We do not assume that the attacker has access to the internal states of LM, including model parameters (Yang et al., 2024), logits for all tokens in the response, and any additional APIs like logit bias that could aid in inferring this state (Carlini et al., 2024b). This assumption aligns with the typical deployment of LLMs.
3 Robust Leakage Identification
@@
Hardness of quantifying partial leakage.   Naturally, the system prompt is fully leaked when the attacker’s guess $\bm{g}$ includes the prompt $\bm{p}$ verbatim. However, quantifying partial leakage in more realistic scenarios–such as when $\bm{g}$ includes a modified version of $\bm{p}$–is challenging. This difficulty stems from two primary factors. First, defining what constitutes private information within $\bm{p}$ is inherently complex. Even if a clear definition is established, the leakage of this information tends to be context-specific and is hard to quantify by comparing $\bm{g}$ and $\bm{p}$ in their utterance (e.g., with BLEU (Papineni et al., 2002) or ROUGE-L scores (Lin, 2004)) or their semantics (e.g., with cosine similarity between text embeddings). Second, $\bm{g}$ may not represent the optimal guess the attacker can make, meaning any insights derived from $\bm{g}$ could underestimate the true extent of the leakage.
This motivates us to consider the worst-case scenario, where leakage occurs if the response $\bm{r}$ the attacker observes [^2] For simplicity, we hereafter assume the attacker makes $k=1$ query unless otherwise stated. contains any information about the system prompt $\bm{p}$. This accounts for the extreme case where the entire $\bm{p}$ is sensitive and for the most powerful attacker capable of losslessly extracting all the information about $\bm{p}$ from $\bm{r}$. With this criterion in mind, a defense is considered effective when $\bm{r}$ reveals no information about $\bm{p}$, or formally $I(\bm{r};\bm{p})=0$, where $I(\mathrm{X};\mathrm{Y})$ represents the mutual information between random variables $\mathrm{X}$ and $\mathrm{Y}$.
Hypothesis testing for zero leakage.   The question of distinguishing zero leakage from other scenarios naturally leads to hypothesis testing, a widely used approach (Kairouz et al., 2015; Nasr et al., 2023). In this context, the null hypothesis $H_{0}$ and alternative one $H_{1}$ are defined as follows:
$H_{0}$ $:\ I(\bm{r};\bm{p})>0,$ (1)
$H_{1}$ $:\ I(\bm{r};\bm{p})=0.$
To perform it, it is natural to consider two distributions: $Q_{\textrm{zero}}(\bm{p},\bm{q})$, the distribution of responses $\bm{r}$ conditioned on $I(\bm{r};\bm{p})=0$, and the counterpart $Q_{\textrm{other}}(\bm{p},\bm{q})$, i.e., the distribution of responses r conditioned on $I(\bm{r};\bm{p})>0$. Denoting the probability density functions for them as $f_{\bm{p},\bm{q}}^{\textrm{zero}}(\cdot)$ and $f_{\bm{p},\bm{q}}^{\textrm{other}}(\cdot)$, respectively, one can define the likelihood ratio $\mathrm{\Lambda}$ as follows:
$\mathrm{\Lambda}(\bm{r};\bm{p},\bm{q})=f_{\bm{p},\bm{q}}^{\textrm{other}}(\bm{r})/f_{\bm{p},\bm{q}}^{\textrm{zero}}(\bm{r}).$ (2)
@@
According to the Neyman Pearson lemma (Neyman & Pearson, 1933), for a target false positive rate $\alpha$, the highest true positive rate $\beta$ among all possible tests is achieved by rejecting $H_{0}$ when $\mathrm{\Lambda}<c$, where $c$ is chosen such that $\Pr[\mathrm{\Lambda}<c\mid H_{0}]=\alpha$ [^3] A false positive occurs when the test incorrectly indicates zero leakage when leakage actually exists, while a true positive indicates correctly detected non-zero leakage. Unfortunately, the multivariate distributions $Q_{\textrm{zero}}$ and $Q_{\textrm{other}}$ lack closed-form expressions, making their direct evaluation challenging. To address this, we propose to approximate them as $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$, the distributions of the mean log-likelihood of model responses conditioned on $I(\bm{r};\bm{p})=0$ and $I(\bm{r};\bm{p})>0$, respectively, where the mean log-likelihood $\mathrm{M}$ of $\bm{r}$ given $\bm{p}$ and $\bm{q}$ is evaluated over all its tokens $r_{1},\dots,r_{n}$ in the spirit of language modeling:
$\mathrm{M}(\bm{r};\bm{p},\bm{q})=\frac{1}{n-1}\sum_{l=0}^{n-1}\log{\Pr[r_{l+1}\mid\bm{p},\bm{q},r_{1},r_{2},\dots,r_{l}]}.$ (3)
@@
$\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})=g_{\bm{p},\bm{q}}^{\textrm{other}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right)/g_{\bm{p},\bm{q}}^{\textrm{zero}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right).$ (4)
@@
To address this, we note that a response generated with $\bm{p}$ should exhibit statistical dependence on $\bm{p}$, implying mutual information exists between the two. Thus, we approximate $\tilde{Q}_{\textrm{other}}$ with $\tilde{Q}^{*}_{\textrm{other}}$, which represents the distributions of the mean log-likelihood of model responses generated with $\bm{p}$ across all possible real-world queries.
We further assume that the LM inherently contains no mutual information with $\bm{p}$, as otherwise $\bm{p}$ would become redundant. Under this assumption, responses will have no mutual information with $\bm{p}$ if and only if the queries themselves are independent with $\bm{p}$. We thus approximate $\tilde{Q}_{\textrm{zero}}$ with $\tilde{Q}^{*}_{\textrm{zero}}$, which represents the distributions of the mean log-likelihood of model responses generated without $\bm{p}$ across all possible real-world queries that have no mutual information with $\bm{p}$.
These approximations make the offline estimation of $\tilde{Q}^{*}_{\textrm{zero/other}}$ feasible (see Appendix C for implementation details). Drawing on established practices (Leino & Fredrikson, 2020; Carlini et al., 2022), we model $\tilde{Q}^{*}_{\textrm{zero/other}}$ as Gaussians. This reduces the estimation process to determining only two parameters—mean and variance—for each distribution. Consequently, we achieve practical offline estimation with minimal sample requirements and computational effort.
4 Defense via On-Demand Regeneration

### v2

# skipabove=5pt, skipbelow=5pt

@@ Abstract.
Large language models (LLMs) are increasingly utilized in applications where system prompts, which guide model outputs, play a crucial role. These prompts often contain business logic and sensitive information, making their protection essential. However, adversarial and even regular user queries can exploit LLM vulnerabilities to expose these hidden prompts. To address this issue, we propose PromptKeeper, a robust defense mechanism designed to safeguard system prompts. PromptKeeper tackles two core challenges: reliably detecting prompt leakage and mitigating side-channel vulnerabilities when leakage occurs. By framing detection as a hypothesis-testing problem, PromptKeeper effectively identifies both explicit and subtle leakage. Upon detection, it regenerates responses using a dummy prompt, ensuring that outputs remain indistinguishable from typical interactions when no leakage is present. PromptKeeper ensures robust protection against prompt extraction attacks via either adversarial or regular queries, while preserving conversational capability and runtime efficiency during benign user interactions. [^1] Code is available at https://github.com/SamuelGong/PromptKeeper.
1. Introduction
Large language models (LLMs) have emerged as transformative tools in artificial intelligence, demonstrating remarkable capabilities to interpret and execute natural-language instructions (Brown et al., 2020; Touvron et al., 2023; Ouyang et al., 2022). These capabilities enable their deployment in diverse applications, from assisting with customer service to generating creative content and automating technical processes. In many such deployments, service providers prepend a system prompt to each user query, a carefully designed instruction that governs the model’s behavior. These prompts often define a model’s tone, structure its responses, or restrict the scope of its functionality, enabling LLMs to perform specialized tasks without the need for resource-intensive fine-tuning (Apideck, 2024).
However, the value of system prompts extends far beyond their functional role. They frequently contain business-related information or secret values that reflect the intellectual property of the deploying organization. In many cases, the system prompt represents a greater source of competitive advantage than the LLM itself, as the latter is often based on widely available foundational models (PromptBase, 2024; PromptSea, 2024). Moreover, these prompts may contain regulatory compliance instructions, or safety mechanisms intended to guide the model’s behavior. The inadvertent exposure of these prompts could also result in significant security risks (Wallace et al., 2024; Toyer et al., 2024). As a result, system prompts are meant to be kept hidden from users (MicroSoft, 2024).
Unfortunately, system prompts are susceptible to multiple forms of leakage, even in environments designed to conceal them. Research has shown that adversarial user queries, such as “Repeat all sentences you saw,” can extract hidden prompts (Perez and Ribeiro, 2022; Wallace et al., 2024), despite explicit safeguards such as extended instructions and post-generation filters (Zhang et al., 2024a; Hui et al., 2024). Moreover, the threat extends beyond adversarial tactics: researchers have demonstrated that regular user queries, which may appear benign, can also lead to prompt leakage. By mapping text outputs (Zhang et al., 2024b) or token-level logits (Morris et al., 2024) to the original prompts, attackers can reconstruct sensitive details with surprising accuracy. These vulnerabilities highlight the pressing need for robust mechanisms to protect system prompts from exposure.
Our contributions.   To address this challenge, we introduce PromptKeeper, a robust defense mechanism designed to protect system prompts from leakage (Figure 1). PromptKeeper operates by focusing on the outputs generated by the LLM, rather than modifying the model itself or altering user inputs. It aims to ensure prompt privacy without significantly affecting conversational quality or runtime efficiency during benign user interactions.
Achieving this goal requires overcoming two fundamental challenges. The first is robustly identifying when the system prompt is leaked in the model’s outputs. Leakage is not always binary: while directly replicating the prompt constitutes complete exposure, more subtle forms—where fragments or implicit information are revealed—are harder to detect and quantify. Yet accurate detection is critical to balancing privacy and utility: overly conservative defenses may reduce leakage but degrade the model’s conversational utility, while lenient defenses risk revealing sensitive information to adversaries. PromptKeeper tackles this by formulating leakage identification as a hypothesis-testing problem. By modeling outputs generated with and without the system prompt, PromptKeeper detects deviations that suggest prompt-related information is leaked in the response. This statistical approach provides a robust and tunable way to identify leakage, without relying on imperfect or fixed metrics such as BLEU (Papineni et al., 2002) or ROUGE-L (Lin, 2004) (Section 3).
Once leakage is detected, the second challenge is determining how to return a response that protects the system prompt while mitigating side-channel privacy vulnerabilities. A naive approach might deny the request or produce a generic error message when leakage is identified, but such behavior creates side channels that attackers can exploit to infer prompt details through patterns in denials. To counter this, PromptKeeper adopts a new response-regeneration strategy. When prompt leakage is detected, it regenerates a new response using a dummy prompt. This dummy prompt mirrors the original prompt’s structure but contains only general, non-sensitive instructions. This ensures that the regenerated response is indistinguishable from typical outputs produced when no leakage occurs, thereby neutralizing adversarial attempts to extract the prompt. Furthermore, because PromptKeeper regenerates responses only when necessary, it preserves both the model’s computational efficiency and conversational utility during benign interactions (Section 4).
@@ 1. Introduction
We evaluate PromptKeeper to assess its effectiveness in safeguarding various system prompts, including those used in real-world GPT Store apps (Section 5). The evaluation encompasses system prompt extraction attacks conducted through both adversarial and regular user queries. Moreover, we introduce a novel method to quantify the protected model’s conversational capability, focusing on its ability to adhere to the scope and behavior defined by the system prompt during benign user interactions. Our extensive experiments demonstrate that PromptKeeper successfully balances system prompt privacy with the model’s adherence to its intended behavior across different LLMs, establishing it as a comprehensive solution in modern LLM-powered systems (Section 6).
@@ 2. Threat Model
Scenario.   As studied in a related work (Zhang et al., 2024a), we consider a scenario where a service API, denoted as $f_{\bm{p}}$, provides text-generation capabilities. The API takes as input a user query $\bm{q}$ and passes to a language model LM, which generates a response $\bm{r}\leftarrow\texttt{LM}(\bm{p},\bm{q})$ using a system prompt $\bm{p}$ secretly owned by the service provider, as well as some employed randomness. In practice, end users may interact directly with $f_{\bm{p}}$, or indirectly via popular application interfaces (OpenAI, 2024b). Depending on the system’s design (e.g., GPT-4 (Wallace et al., 2024) vs. GPT-3 (Mann et al., 2020)), $\bm{p}$ and $\bm{q}$ may be processed separately with different privilege levels, or simply concatenated before being fed to LM. Regardless of the implementation, the essential point is that the system prompt is never directly revealed to the user; it is intended to remain confidential while influencing the generated response in aspects including but not limited to roles, styles, and contexts.
System prompt extraction.   The attacker’s goal is to accurately guess the system prompt $\bm{p}$ by using a set of responses $\bm{r}_{1},\dots,\bm{r}_{k}$ acquired through $k$ queries made to the API using $\bm{q}_{1},\dots,\bm{q}_{k}$. The guess $\bm{g}$ is generated as $\bm{g}=\texttt{recon}(\bm{r}_{1},\dots,\bm{r}_{k})$, where $\texttt{recon}(\cdot)$ denotes any reconstruction mechanism the attacker wishes to use. Such mechanisms could be as simple as direct string manipulation or as complex as using a large neural network to infer patterns from the observed responses. As for the attacker’s capabilities, we further make the following assumptions:
  1. (1)
    Black-box access only: The attacker interacts with the service solely via typical public APIs. They cannot inspect the model parameters (weights), internal states (LM hidden layers), or token-level logits (Yang et al., 2024).
  2. (2)
    Limited side-information: The attacker does not have additional privileged data, such as logit bias values, system logs, or side-channel measurements that could aid in inferring $\bm{p}$ (Carlini et al., 2024b).
  3. (3)
    Adaptive querying: The attacker can adapt their queries based on previous responses. By observing how $\bm{r}$ changes with $\bm{q}$, they may attempt to coax the LM into revealing hints about $\bm{p}$.
  4. (4)
    Knowledge of LLM architecture (optional): In some cases, the attacker may know the approximate model architecture (e.g., GPT-like), publicly available release notes, or the training corpus. However, we do not assume they know exact internal hyperparameters for training, or the specific system prompt design choices.
These assumptions align with the typical deployment of LLMs.
3. Robust Leakage Identification
Prompt privacy vs. prompt adherence.   According to information theory, the only way to ensure perfect privacy for the system prompt, $\bm{p}$, is by not providing it to the model at all. However, this approach eliminates prompt adherence—the ability of the model to follow specific requirements, guidelines, or constraints encoded in $\bm{p}$—nullifying the purpose of a carefully crafted system prompt. Conversely, if one employs no protections against system prompt disclosure, she could enjoy full adherence to the prompt but risk exposing $\bm{p}$ entirely. In practice, achieving a balance between preserving the confidentiality of $\bm{p}$ and ensuring its influence on the model’s outputs presents a critical tradeoff.
Challenges in quantifying partial leakage.   Balancing privacy and adherence involves regulating how much of $\bm{p}$ is revealed, either directly or indirectly, through the model’s outputs. Quantifying this “leakage” is complex. Apparently, a system prompt is fully leaked when an attacker’s guess, $\bm{g}$, includes the prompt $\bm{p}$ verbatim. However, measuring partial leakage in more realistic scenarios–such as when $\bm{g}$ includes a modified version of $\bm{p}$–is particularly challenging. This difficulty stems from two primary factors:
  1. (1)
@@ 3. Robust Leakage Identification
    Defining private information: Identifying what constitutes private information within $\bm{p}$ is inherently complex. Even if a clear definition is established, the leakage of this information tends to be context-specific and is hard to quantify by comparing $\bm{g}$ and $\bm{p}$ in their utterance (e.g., with BLEU (Papineni et al., 2002) or ROUGE-L scores (Lin, 2004)) or their semantics (e.g., with cosine similarity between text embeddings).
  2. (2)
    Suboptimal attacker guesses: An attacker’s guess $\bm{g}$ may not represent the optimal guess the attacker can make, meaning any insights derived from $\bm{g}$ could underestimate the true extent of the leakage.
While the second issue can be approached by examining the attacker’s input (i.e., the model’s response $\bm{r}$) rather than $\bm{g}$, solving the first issue requires a principled approach to measure or bound leakage, avoiding unreliable ad-hoc proxies.
Zero leakage as a first-principle baseline.   In the absence of a reliable metric for partial leakage, we use zero leakage as a baseline for evaluation. Specifically, we first ask: if no prompt $\bm{p}$ were used (implying no leakage), how would the model’s outputs be distributed? For any actual response $\bm{r}$ generated using $\bm{p}$, we then assess how likely it is to arise from this “zero-leakage” scenario. This approach naturally lends itself to a hypothesis testing framework, a widely used method in the privacy literature to distinguish between competing scenarios (Kairouz et al., 2015; Nasr et al., 2023). In this context, the null hypothesis $H_{0}$ and alternative hypothesis $H_{1}$ are defined as follows:
(1) $H_{0}$ $:\ I(\bm{r};\bm{p})>0,$
$H_{1}$ $:\ I(\bm{r};\bm{p})=0,$
where $I(\mathrm{X};\mathrm{Y})$ represents the mutual information between random variables $\mathrm{X}$ and $\mathrm{Y}$. Although $H_{1}$ (zero leakage) is not a practical operating point—since using $\bm{p}$ always introduces some dependence—it functions as an anchor for a full-spectrum assessment. By design, we test how “prompt-like” a response appears to be compared with a scenario in which $\bm{p}$ was never employed.
Hypothesis testing with a tunable tolerance.   We operationalize this baseline through likelihood ratio tests, comparing the likelihood of observing $\bm{r}$ under two distributions: $Q_{\textrm{zero}}$ (for the zero-leakage world) and $Q_{\textrm{other}}$ (for the non-zero leakage world). Denoting their probability density functions for them as $f_{\bm{p},\bm{q}}^{\textrm{zero}}(\cdot)$ and $f_{\bm{p},\bm{q}}^{\textrm{other}}(\cdot)$, respectively, the likelihood ratio $\mathrm{\Lambda}$ is defined as:
(2) $\mathrm{\Lambda}(\bm{r};\bm{p},\bm{q})=f_{\bm{p},\bm{q}}^{\textrm{other}}(\bm{r})/f_{\bm{p},\bm{q}}^{\textrm{zero}}(\bm{r}).$
@@
According to the Neyman Pearson lemma (Neyman and Pearson, 1933), for a target false positive rate $\alpha$, the highest true positive rate $\beta$ among all possible tests is achieved by rejecting $H_{0}$ when $\mathrm{\Lambda}<c$, where $c$ is chosen such that $\Pr[\mathrm{\Lambda}<c\mid H_{0}]=\alpha$ [^2] A false positive occurs when the test incorrectly indicates zero leakage when leakage actually exists, while a true positive indicates correctly detected non-zero leakage.
In practice, both $Q_{\textrm{zero}}$ and $Q_{\textrm{other}}$ are high-dimensional, and their closed-form expressions are not readily available. To bridge theory and practice, we propose to approximate them as $\tilde{Q}_{\textrm{zero}}(\bm{p},\bm{q})$ and $\tilde{Q}_{\textrm{other}}(\bm{p},\bm{q})$, the distributions of the mean log-likelihood of model responses conditioned on $I(\bm{r};\bm{p})=0$ and $I(\bm{r};\bm{p})>0$, respectively, where the mean log-likelihood $\mathrm{M}$ of $\bm{r}$ given $\bm{p}$ and $\bm{q}$ is evaluated over all its tokens $r_{1},\dots,r_{n}$ in the spirit of language modeling:
(3) $\mathrm{M}(\bm{r};\bm{p},\bm{q})=\frac{1}{n-1}\sum_{l=0}^{n-1}\log{\Pr[r_{l+1}\mid\bm{p},\bm{q},r_{1},r_{2},\dots,r_{l}]}.$
@@
(4) $\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})=g_{\bm{p},\bm{q}}^{\textrm{other}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right)/g_{\bm{p},\bm{q}}^{\textrm{zero}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right).$
In essence, evaluating leakage boils down to checking whether $\mathrm{M}(\bm{r};\bm{p},\bm{q})$ aligns more with the “zero leakage” fit or the “non-zero leakage” fit. The hyperparameter $\alpha$ can be deemed as the tolerance level for tuning how aggressively we flag suspicious responses for disclosing too much about $\bm{p}$. It is also worth mentioning that our approach parallels differential privacy (DP) (Dwork and Roth, 2014; Cynthia, 2006), where perfect zero leakage aligns with $\epsilon=0$. Similar to how DP tolerates $\epsilon>0$ to balance utility and privacy, we allow a tunable significance level $\alpha$ to manage the tradeoff between prompt privacy and adherence. This analogy is not coincidental—the relationship between privacy leakage identification as hypothesis testing and differential privacy has been previously explored, e.g., in Kairouz et al. (2015).
Figure 2. Standard text generation workflow with major defenses for system prompt privacy.
@@
To address this, we note that a response generated with $\bm{p}$ should exhibit statistical dependence on $\bm{p}$, implying mutual information exists between the two. Thus, we approximate $\tilde{Q}_{\textrm{other}}$ with $\tilde{Q}^{*}_{\textrm{other}}$, which represents the distributions of the mean log-likelihood of model responses generated with $\bm{p}$ across all possible real-world queries. We further assume that the LM inherently contains no mutual information with $\bm{p}$, as otherwise $\bm{p}$ would become redundant. Under this assumption, responses will have no mutual information with $\bm{p}$ if and only if the queries themselves are independent with $\bm{p}$. We thus approximate $\tilde{Q}_{\textrm{zero}}$ with $\tilde{Q}^{*}_{\textrm{zero}}$, which represents the distributions of the mean log-likelihood of model responses generated without $\bm{p}$ across all possible real-world queries that have no mutual information with $\bm{p}$.
These approximations make the offline estimation of $\tilde{Q}^{*}_{\textrm{zero/other}}$ feasible (see Section 5.3 for implementation details). Drawing on established practices (Leino and Fredrikson, 2020; Carlini et al., 2022), we model $\tilde{Q}^{*}_{\textrm{zero/other}}$ as Gaussians. This reduces the estimation process to determining only two parameters—mean and variance—for each distribution. Consequently, we achieve practical offline estimation with minimal sample requirements and computational effort.
Putting it all together.   To facilitate the reader’s understanding of our proposed leakage identification approach, we briefly summarize the end-to-end workflow below:
  1. (1)
    For a response $\bm{r}$ under evaluation, its mean log-likelihood $\mathrm{M}(\bm{r};\bm{p},\bm{q})$ is obtained as a by-product of the generation process.
  2. (2)
    Using the distributions $\tilde{Q}^{*}_{\textrm{zero}}$ and $\tilde{Q}^{*}_{\textrm{other}}$ pre-computed offline, compute the two probability densities $g_{\bm{p},\bm{q}}^{\textrm{zero}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right)$ and $g_{\bm{p},\bm{q}}^{\textrm{other}}\left(\mathrm{M}(\bm{r};\bm{p},\bm{q})\right)$ for the obtained mean log-likelihood value, respectively.
  3. (3)
    Compute the approximated likelihood ratio $\mathrm{\tilde{\Lambda}}(\bm{r};\bm{p},\bm{q})$ based on these two densities to perform hypothesis testing at a predefined significance level $\alpha$.
Figure 3. Example of the side-channel created by denial of service during response-based protection.
4. Defense via On-Demand Regeneration
4.1. Taxonomy of Existing Defense Schemes